### v1

# OBI-Bench: Can LMMs Aid in Study of Ancient Script on Oracle Bones?

Abstract
We introduce OBI-Bench, a holistic benchmark crafted to systematically evaluate large multi-modal models (LMMs) on whole-process oracle bone inscriptions (OBI) processing tasks demanding expert-level domain knowledge and deliberate cognition. OBI-Bench includes 5,523 meticulously collected diverse-sourced images, covering five key domain problems: recognition, rejoining, classification, retrieval, and deciphering. These images span centuries of archaeological findings and years of research by front-line scholars, comprising multi-stage font appearances from excavation to synthesis, such as original oracle bone, inked rubbings, oracle bone fragments, cropped single character, and handprinted character. Unlike existing benchmarks, OBI-Bench focuses on advanced visual perception and reasoning with OBI-specific knowledge, challenging LMMs to perform tasks akin to those faced by experts. The evaluation of 6 proprietary LMMs as well as 17 open-source LMMs highlights the substantial challenges and demands posed by OBI-Bench. Even the latest versions of GPT-4o, Gemini 1.5 Pro, and Qwen-VL-Max are still far from public-level humans in some fine-grained perception tasks. However, they perform at a level comparable to untrained humans in deciphering task, indicating remarkable capabilities in offering new interpretative perspectives and generating creative guesses. We hope OBI-Bench can facilitate the community to develop domain-specific multi-modal foundation models towards ancient language research and delve deeper to discover and enhance these untapped potentials of LMMs.
Figure 1: Overview of the OBI-Bench. OBI-Bench presents five in-process tasks: 1) recognition: locating dense oracle bone characters from original oracle bone or rubbings; 2) rejoining: reconstructing fragmented text fragments into coherent texts; 3) classification: categorizing individual characters into their respective meanings; 4) retrieval: returning relevant results according to the given query OBI images; 5) deciphering: interpreting the OBI for historical and cultural investigation.
1 Introduction
Oracle bone inscriptions (OBI) have been recognized as valuable records of making divination and praying to gods by the late Shang people from 1400 B.C. to 1100 B.C., in addition to being the earliest evidence of a Chinese writing system (UNESCO, 2017). Their discovery and interpretation provide a unique window for evidence of the thought, language, society, and history of past civilizations. All along, there have been several challenges from the excavation to the interpretation of OBI. First, after thousands of years buried in the ground, the original OBIs are inherently eroded, fragmented, and susceptible to damage. Recognizing oracle bone characters from scattered pieces not only helps in initially determining the existence of the characters but also aids in their digital archiving with minimal physical handling. Second, the rejoining work of OBI is an indispensable prerequisite for follow-up research, which restores the original appearance of oracle Bones and offers complete and accurate information for OBI scholars (Zhang et al., 2022). Current rejoining procedures are non-trivial and involve time-consuming and specialized workflows (Men, 2008), thus posing demands to develop automatic or AI-assisted rejoining techniques. Third, OBIs exhibit a wide range of stylistic variations in font and normally suffer from various types of noise in real rubbings (e.g., stroke broken, edge-cracked, and spindles), making it arduous to distinguish from distortions and accurately classify them. Fourth, OBI retrieval, as a core component for building large-scale databases, its difficulty lies in differentiating the inter-class and intra-class discrepancies of oracle bone characters. Fifth, among the thousands of OBIs that have been excavated, over half of them remain undeciphered (Wang & Deng, 2024). The traditional methods for oracle character deciphering typically involve a combination of historical linguistics and archaeological context. More recently, lots of efforts (Jiang et al., 2023; Gao et al., 2024; Qiao et al., 2024; Guan et al., 2024b) have been made to leverage the representation capability of neural networks so as to establish correspondences between OBIs and modern Chinese. However, many characters have evolved in meaning through cultural and historical changes, and not all oracle bone characters can be translated into modern Chinese with synonymous expressions. Another problem lies in the absence of a comprehensive corpus and the lack of OBI-encoding systems, which limits the applicability of language models. These five tasks are crucial steps towards placing an OBI both in history and within the world of the people who read and study it.
Nowadays, the emergent large multi-modal models (LMMs) have brought opportunities for solving multidisciplinary tasks with powerful visual perception, understanding, and reasoning abilities (Yue et al., 2024; Caffagni et al., 2024; OpenAI, 2024b). Meanwhile, the ability of LMMs remains unclear on fine-grained perception and cognition (i.e., high-level perception), which play significant roles in interpreting ancient texts and its associated tasks on image material pre-processing. The perception at different granularities is strongly associated with a wide range of OBI tasks, such as distinguishing noise from real oracle bone character recognition (Wang et al., 2022), marginal sealability checking in rejoining (Zhang et al., 2021c) and component-level structural decomposition in retrieval (Hu et al., 2024). Henceforth, it is worth evaluating the current abilities of LMMs in fine-grained perception and cognition based on actual needs to relieve extensive human resources and seek feasible solutions for future OBI research.
In this work, we propose the first comprehensive OBI-focused benchmark, OBI-Bench, to evaluate the recent LMMs in whole-process OBI tasks including recognition, rejoining, classification, retrieval, and deciphering under different settings. Our benchmark is constructed around a key question:
Can LMMs aid in study of ancient script on oracle bones?
Specifically, we define two general evaluation principles for LMMs in targeting OBI problems:
  • Task-oriented Perception. As shown in the bottom of Fig. 1, an LMM is expected to answer accurately to achieve the objectives of five in-process tasks like an OBI specialist, such as bounding the position of each oracle bone character for the recognition task or providing a reasonable interpretation of OBIs for the deciphering task.
  • Spanning from excavation to synthesis. As shown in the middle of Fig. 1, an LMM should be able to handle five different forms of oracle bone inscriptions and show good adaptiveness to the appearance or structural variants of OBI within the same task.
To systematically evaluate the whole-process perception ability on various visual granularity under diverse OBI tasks, we collect 5,523 OBI images from 11 distinct sources. Due to the lack of publicly available OBI recognition datasets on real oracle bones and OBI rejoining datasets, we propose the original oracle bone recognition (O2BR) dataset and OBI-rejoin dataset as complements. Aligned with existing practices (Yue et al., 2024; Wu et al., 2024), each image in OBI-Bench is equipped with a question alongside a correct answer. Moreover, we craft a spectrum of questions: What question, Yes-or-No question, How question, and Where question, evaluating from coarse-grained perception to finer-grained perception. A total of 23 popular LMMs are selected for evaluation including 6 proprietary LMMs (e.g., GPT-4o, Qwen-VL-Max, and Gemini 1.5 Pro) and 17 open-source LMMs (e.g., LLaVA-Next, InternVL2, and mPLUG-Owl3). We showcase some empirical findings here: 1) LMMs still have much room for improvement in fine-grained OBI recognition tasks such as quantity detection or locating; 2) LMMs are not sensitive enough to local information such as the borders of the fragmented margin to directly meet the requirements of rejoining. Meanwhile, part of LMMs can help OBI scholars narrow down the range of qualified fragments (GPT-4o and Gemini 1.5 Pro achieves an average performance of 77.57% on the Acc@10 metric); 3) LMMs can be applied in classification and retrieval tasks, especially for cleaner handprinted datasets; 4) partial LLMs rival the performance of public-level humans in deciphering tasks and exhibit remarkable visual reasoning ability in interpreting those genuine undeciphered oracle bone characters. The full list of our findings from different OBI tasks is in the Sec. 3.
By evaluating up-to-date LMMs from different perspectives (OBI research demands), we gain insights into their strengths, limitations, and potential directions for improvement. Ultimately, our objective is to facilitate the field of paleography by fostering the development of more reliable, unbiased, and task-oriented LMMs that meet the needs of OBI experts while upholding trustworthy standards and contributing to historical debates.
Table 1: Overview of the 11 different image source datasets in the OBI-Bench, and the respective dataset attributes among recognition, rejoining, classification, retrieval, and deciphering tasks.
Task Image Source Dataset Type Sampled Size Avg. Res. Remarks
Recognition YinQiWenYuandetection (AYNU, 2020) Fragments (inked) 2,000 435$\times$538 Coordinate
O2BR (Ours) Fragments (original) 800 1529$\times$1192 Coordinate
[…]
Oracle-50K (Han et al., 2020) Handprinted 500 50$\times$50 Class label
OBI125 (Yue et al., 2022) Cropped words 500 77$\times$135 Class label
[…]
OBI-IJDH (Fujikawa & Meng, 2020) Cropped words 100 60$\times$91 Class label
[…]
OBI Component 20 (Hu et al., 2024) Handprinted 50 66$\times$64 Interpretation
HUST-OBS (Wang et al., 2024a) Handprinted 40 112$\times$172 Interpretation
2 Constructing the OBI-Bench
2.1 General Principles
Focusing on OBI Task-oriented Abilities of LMMs. Different from existing LMM benchmarks (Liu et al., 2023b; Li et al., 2024b; Liu et al., 2024b; Yue et al., 2024) that either aim at all-round abilities or generic scenarios, the evaluation principles in OBI-Bench are OBI task-oriented. Specifically, we focus on five major issues in the field of oracle bone inscription research: 1) Recognition, involves the positioning and identification of OBI characters on different carriers. 2) Rejoining, refers to the process of stitching broken OBI fragments together to rebuild complete inscriptions. 3) Classification, means categorizing the recognized OBI characters into groups based on their shapes or meanings. 4) Retrieval, is to find a cognate character from large collections of OBIs. 5) Deciphering, includes interpreting the meaning of the oracle bone characters even the contextual semantic information. We adhere to the principles in designing the visual and cognitive tasks, making the proposed OBI-Bench a focused reflection on the OBI processing abilities of LMMs.
Covering Multi-stage Font Appearances. To cover diverse appearances of the oracle bones since their excavation, we collect multi-sourced images for each task, as depicted in Tab. 1. In particular, considering that there are currently no available original oracle bone recognition datasets and rejoining datasets, we proposed the O2BR and OBI-rejoin datasets respectively (See more details in App. B). Due to the specific requirement of domain knowledge in OBI research, three domain experts and one senior OBI scholar were involved in the annotation process of O2BR and OBI-rejoin as well as the sampling process of all source datasets. We include all the manifestations of the entire process of OBI processing from excavation (i.e., original oracle bone) to artificial synthesis (i.e., handprinted or computer-generated). The diverse and multiple sources of images morph the OBI-Bench into a holistic and balanced benchmark to fairly evaluate the competency of LMMs in OBI-related tasks.
2.2 Benchmark from Visual Ability to Cognitive Ability
In the five tasks of OBI-Bench, we evaluate the visual perception and high-level cognitive abilities of LMMs to examine whether they can handle OBI problems by using simple natural queries. For this purpose, we collect 5,523 images (I) from multiple sources (Tab. 1) with diverse task concerns. Then, we design different question prompts (Q) based on the type of tasks and obtain the correct answer (A) from the corresponding dataset for each image. The 5,523 (I, Q, A) tuples compose into the first visual question answering (VQA) testbench (Fig. 2) in the OBI field. Specifically, the questions in OBI-Bench cover two quadrants of concerns for evaluation (Sec. 2.2.1) and four question types (Sec. 2.2.2). The details are elaborated as follows.
Figure 2: Sampled OBI-Bench examples from each task. 5,523 (I,Q,A) tuples span two quadrants of OBI concerns and encompass four types of questions, providing an all-around evaluation of the ability of LMMs on OBI tasks. Note that OBI classification and retrieval tasks share the same queries.
2.2.1 Quadrants for Evaluation Concerns
Axis 1: From Coarse-grained Perception to Fine-grained Perception. The primary axis differentiates the visual perception granularity of OBI tasks. First, coarse-grained perception focuses on the basic forms, themes, or overall structures of the oracle bone inscriptions, which distinguishes OBIs from other symbols or ancient characters (Qiao et al., 2024). Second, fine-grained perception (Xing et al., 2019; Fujikawa et al., 2023) not only involves identifying subtle characteristics, quantity, and spatial positions of OBIs but also reflects the high-level cognitive ability of LMMs (e.g., OBI domain expertise and interdisciplinary reasoning). Several studies (Liu et al., 2023a; He et al., 2024; Tong et al., 2024) followed this paradigm and built on it to refine pixel-level perception while extending interactions beyond text-specific inputs.
Axis 2: From Excavation to Synthesis. In the era predating the prevalence of artificial intelligence (AI), unearthed oracle bone fragments were normally processed through the manual rubbing and image cropping processes for preservation. After thousands of years of natural weathering, corrosion, and man-made destruction, there are various types of noise in oracle bone rubbings. To improve the readability of the original OBIs and expand the data size that could be processed by deep learning algorithms, researcher obtained pseudo-oracle bone characters through handwriting or generative models (Han et al., 2020; Li et al., 2020; Wang et al., 2024b). Acknowledging these forms of OBIs, we curate 5,523 OBI images from different stages of processing, that require LMMs to grasp the content or other relevant domain knowledge to answer correctly.
2.2.2 Question Types
In OBI-Bench, we design four question types, What, Yes-or-No, How, and Where, to simulate multiple human queries in various OBI tasks, each tailored to different levels of granularity as follows.
$\bullet$ What Question. The What questions are a common type of query in some LMM benchmarks (Wu et al., 2024; Zhang et al., 2024b). In OBI-Bench, they serve as global coarse-grained perception in recognition task (e.g., What is in this image?), or unbound the answers in deciphering task (e.g., What is the meaning of this oracle bone character?). In contrast to other question types, the What questions examine more comprehensive visual understanding and cognitive abilities of LMMs, by requiring correct attribute perception and knowledge integration.
$\bullet$ Yes-or-No Question. As a fundamental type of judgment, Yes-or-No represents a binary output. This design minimizes the ambiguity of the model answers and is closer to the task objectives than the open-ended responses of What questions.
$\bullet$ How Question. Except from two common types, we also include the How questions to further refine the responses as an extension to Yes-or-No questions. As shown in Fig. 2, we can query How many oracle bone characters are in this image? for the recognition task (Fu et al., 2023) or query How much probability is there that these two oracle bone characters belong to the same class? for the classification task to derive quantitative results predicted by LMMs.
$\bullet$ Where Question. We specifically employ the Where question for the recognition task. For instance, we can query Return a bounding box for each detected oracle bone character in [xmin, ymin, xmax, ymax] format to achieve pixel-level OBI anchoring, which requires the LMM to have a fine-grained perception, including spatial location, shape, and relationships between detected objects.
3 Experiments
In OBI-Bench, we evaluate 17 up-to-date popular and competitive open-source LMMs, together with 6 proprietary LMMs, under zero-shot settings. More results and analyses are appended in App. C.
Table 2: Performance comparison on the OBI recognition task. The best result is marked in bold and the second-best result is underlined for both proprietary and open-source LMMs respectively.
Datasets O2BR YinQiWenYuandetection
[…]
Human (public) 0.9364 100% 0.0033 0.9272 0.9189 100% 0.0060 0.8776
Proprietary LMMs:
[…]
Gemini 1.5 Flash 0.4123 96.50% 1.7857 0.0962 0.3425 97.35% 0.4522 0.0276
GPT-4v 0.5408 99.30% 0.4223 0.0022 0.3701 99.65% 0.4383 0.0165
GPT-4o (ver. 0806) 0.6114 99.95% 0.4016 0.0038 0.3734 99.90% 0.3458 0.0182
Qwen-VL-Max (ver. 0809) 0.6071 99.63% 0.4799 0.0086 0.3375 99.55% 0.4843 0.0131
GLM-4V 0.5319 39.17% 0.3681 0.0041 0.3635 52.45% 0.3632 0.0124
Open-source LMMs:
[…]
mPLUG-Owl3 (Qwen2-7B) 0.3342 99.88% 0.4474 0.0811 0.2505 99.95% 0.6593 0.0527
MiniCPM-V 2.6 (Qwen2-7B) 0.5576 88.75% 0.4829 0.0384 0.3781 99.10% 1.1793 0.0111
Moondream2 (ver. 0728) 0.4818 98.25% 0.6795 0.0400 0.3049 91.30% 0.4436 0.0653
InternVL2-Llama3-76B (Llama3-70B) 0.5833 99.65% 0.4328 0.0976 0.3892 99.75% 0.5344 0.0623
InternVL2-40B (Nous-Hermes2-Yi-34B) 0.5664 98.35% 0.4561 0.0766 0.3637 99.05% 0.6733 0.0487
InternVL2-8B (InternLM2.5-7B) 0.5232 95.00% 0.4618 0.0020 0.3429 97.95% 1.1146 0.0152
GLM-4V-9B (GLM-4-9B) 0.5388 29.50% 0.4825 <10e-4 0.2839 15.70% 0.3934 <10e-4
CogVLM2-Llama3-19B (Llama3-8B) 0.5321 61.00% 0.6928 <10e-4 0.3966 91.75% 0.5568 0.0002
LLaVA-NeXT (Qwen1.5-72B) 0.4846 99.75% 1.1011 0.0445 0.3297 99.75% 1.0561 0.0591
LLaVA-NeXT (Llama3-8B) 0.4764 93.13% 0.5512 0.0001 0.3120 93.90% 0.4268 0.0189
IDEFICS-2-8B (Mistral-7B) 0.3175 95.88% 0.4916 <10e-4 0.2658 95.10% 0.5119 <10e-4
DeepSeek-VL (DeepSeek-LLM-7B) 0.5111 92.75% 0.5657 0.0449 0.3386 98.00% 0.6263 0.0520
InternLM-XComposer2-VL (InternLM2-7B) 0.5106 99.88% 0.6641 0.0049 0.2661 99.95% 1.5231 0.0281
LLaVA-v1.5 (Vicuna-v1.5-13B) 0.4416 99.00% 2.7553 0.0751 0.2875 98.05% 1.7662 0.0493
LLaVA-v1.5 (Vicuna-v1.5-7B) 0.4239 91.88% 0.8861 0.0656 0.2729 81.90% 1.9621 0.0465
Qwen-VL (Qwen-7B) 0.4489 74.13% 1.7416 0.0003 0.3137 86.81% 3.5694 0.0069
3.1 Evaluation on Recognition
In this section, we examine the recognition ability of LMMs, comparing them from coarse-grained global content perception to fine-grained OBI locating.
Setup. To evaluate how well LMMs perform under various OBI recognition scenarios, we design four evaluation schemes across two image sources (i.e., original oracle bone from our self-curated O2BR dataset and inked rubbings from the YinQiWenYuandetection (AYNU, 2020) dataset): (1) evaluating via the What query to measure the preciseness of the global perceptual descriptions on OBIs. Given the commonality of contents in sampled test sets, we predefined reference descriptions for each of the two datasets. The max_tokens for all LMMs is set to $200$. The similarity between text embeddings is used as the evaluation metric. (2) evaluating via the Yes-or-No query after specifying the response direction (the words “OBI” appear explicitly in the query), designed to complement the evaluation on coarse-grained perception. The accuracy of answers is adopted as the evaluation criterion. (3) evaluating via the How query to quantify the ability of LMMs in parsing irregularly structured character-level OBI contents. We use mean relative error (MRE) as the metric. (4) evaluating via the Where query to achieve finer-grained perception (i.e., dense oracle bone character locating), designed to more effectively uncover the practicability of LMMs in real OBI processing pipeline. Mean intersection over union (mIoU) is used as the metric. More details are in App. C.2.
Results. In scenario (1), as shown in Tab. 2, among proprietary LMMs, the recently-released GPT-4o reaches the best performance in terms of the relevance of answers on coarse-grained perception, followed by Qwen-VL-Max, which shows rather close results. By achieving over 12.4% on average performance improvement, these models provide more precise answers than the open-source LMMs. Similar improvement can be observed for the same series of LMM (e.g., LLaVA and InternVL2), as the number of parameters of the language backbone increases. In scenario (2), we add keyword (i.e., oracle bone) directly to a Yes-or-No question to further evaluate the global perception of LMMs. We find most LMMs perform nearly 100% correctness, indicating that these models are highly effective in handling queries that involve explicitly directed content. One exception is the GLM-4V series that only achieves an average accuracy of 34.34%. We speculate it is due to the answer preference bias (Wu et al., 2024). Deeper analysis is in App. C.2.2. In scenario (3), we transition to examining the fine-grained visual perceptual of LMMs, i.e., counting the number of oracle bone characters. Our results in Tab. 2 indicate that GLM-4V and InternVL2-Llama3-76B exhibit the best performance within their respective division groups. Additionally, we observe some abnormally large values (>50) or repeated values from Gemini 1.5 Flash, xGen-MM-4B, LLaVA-NeXT-72B, LLaVA-v1.5-13B, and Qwen-VL that results in their relatively large MRE (Zhou et al., 2024). In scenario (4), all evaluated LMMs consistently underperform. Gemini 1.5 Pro significantly outperforms other LMMs (surpasses GPT-4v and GPT-4o two orders of magnitude) (Ueno & Lynn, 2024) with its exclusive coordinate return function. Overall, current LMMs are still not usable for character-level OBI locating and are far from the public-level human.
Table 3: Results on the OBI rejoining task. ‘Yes-or-No’ and ‘How’ represent the absolute output and probability output, respectively. We report Acc@5 for the ‘Yes-or-No’ query. Given that not all LMMs support multi-image inputs natively, we shrink the number of evaluated models.
[…]
Gemini 1.5 Pro 28.53% 24.88% 46.37% 76.67%
Gemini 1.5 Flash 22.17% 19.33% 36.34% 66.76%
GPT-4v 27.63% 26.86% 43.75% 73.75%
GPT-4o (ver. 0806) 32.21% 29.13% 48.43% 78.47%
Qwen-VL-Max (ver. 0809) 21.77% 13.14% 31.68% 56.67%
GLM-4V 5.33% 4.58% 12.19% 21.46%
[…]
mPLUG-Owl3 (Qwen2-7B) 14.48% 3.52% 12.63% 17.60%
MiniCPM-V 2.6 (Qwen2-7B) 13.23% 2.69% 11.18% 16.98%
InternVL2-Llama3-76B (Llama3-70B) 20.13% 5.18% 16.98% 31.68%
InternVL2-40B (Nous-Hermes2-Yi-34B) 16.00% 3.31% 10.97% 24.84%
InternVL2-8B (InternLM2.5-7B) 14.68% 3.73% 10.14% 17.81%
LLaVA-NeXT (Qwen1.5-72B) 17.45% 4.97% 14.29% 27.74%
LLaVA-NeXT (Llama3-8B) 11.41% 3.93% 9.52% 16.77%
IDEFICS-2-8B (Mistral-7B) 7.38% 3.52% 12.42% 14.91%
DeepSeek-VL (DeepSeek-LLM-7B) 9.40% 4.35% 8.90% 11.59%
Qwen-VL (Qwen-7B) 9.86% 4.02% 8.96% 14.08%
3.2 Evaluation on Rejoining
In this section, we delve into the rejoining performance of LMMs against various eroded, tiny, and irregularly structured fragments of OBI, focusing on the estimation of rejoinable fragments. This task is also a major concern in the recovery of paper money, calligraphy, and painting files recovery.
Setup. Given the absence of publicly available rejoining datasets in OBI domain, we create a dataset consisting of 200 complete OBI pieces across two appearances (i.e., original and inked oracle bone fragments) with 483 rejoinable fragments. More details are in App. B.2. To evaluate the rejoining performance of LMMs on oracle bone fragments with different marginal sealability concerns, we elaborate two evaluation scenarios: (1) evaluating via standard Yes-or-No question that represents the absolute judgments to tentatively explore the rejoining performance; (2) evaluating via How question that outputs rejoinable probabilities for pairs of fragments, allowing for a more accurate and quantitative measurement. Accuracy@k indices are utilized as evaluation metrics.
Results. Tab. 3 presents the evaluation results of LMMs on the OBI rejoining task, from which we gain several valuable observations as follows. First, GPT-4o and InternVL2-Llama3-76B achieve the best performance in their respective categories, although there is still a huge gap between them. Second, among all proprietary LMMs, the rejoining performance under ‘How’ queries is significantly better than that under ‘Yes-or-No’ queries. This indicates that the answer in probabilistic form better reflects the differences in visual perception compared to absolute form output. Third, we notice that GPT-4o and Gemini 1.5 Pro reach over 76% in Acc@10 metric, illustrating the possibility of of using LMMs to assist the OBI rejoining efforts. Fourth, the number of parameters is roughly proportional to the performance, and newer models tend to exhibit superiority under the same scale of language backbone models. Fifth, we find that the overall performance of open-source LMMs is still far away from being truly usable (only 3.85%, 11.39%, and 18.65% on average in terms of Acc@1, Acc@5, and Acc@10, respectively), and their ranks vary greatly under different evaluation criteria.
Table 4: Classification accuracy (%) on HWOBC, Oracle-50k, and OBI125 datasets (delimited by slashes) under different queries. ‘Yes-or-No’ and ‘How’ represent the absolute output and probability output, respectively. We report Acc@5 for the ‘Yes-or-No’ query.
[…]
Gemini 1.5 Pro 68.75%/68.00%/58.50% 88.75%/85.50%/77.50% 100.0%/100.0%/93.75%
Gemini 1.5 Flash 62.50%/60.50%/50.25% 82.00%/82.50%/71.75% 100.0%/99.50%/91.00%
GPT-4v 69.50%/66.00%/57.75% 86.75%/88.25%/70.75% 100.0%/100.0%/91.75%
GPT-4o (ver. 0806) 72.75%/74.50%/62.50% 89.75%/90.25%/75.50% 100.0%/100.0%/93.50%
Qwen-VL-Max (ver. 0809) 64.25%/65.75%/55.00% 85.00%/88.75%/69.25% 100.0%/98.75%/89.75%
GLM-4V 35.50%/29.75%/34.50% 71.00%/68.50%/54.25% 83.50%/85.75%/77.75%
[…]
mPLUG-Owl3 (Qwen2-7B) 39.25%/39.75%/38.75% 44.75%/46.25%/43.75% 56.25%/52.50%/53.50%
MiniCPM-V 2.6 (Qwen2-7B) 40.75%/42.50%/38.50% 45.75%/45.00%/42.75% 56.75%/55.75%/51.75%
InternVL2-Llama3-76B (Llama3-70B) 44.75%/47.50%/43.25% 53.75%/55.00%/50.75% 69.75%/69.00%/66.75%
InternVL2-40B (Nous-Hermes2-Yi-34B) 42.75%/43.50%/40.25% 49.75%/50.00%/48.75% 63.75%/62.25%/59.75%
InternVL2-8B (InternLM2.5-7B) 42.25%/41.75%/38.75% 47.75%/49.00%/47.75% 59.75%/59.00%/56.50%
LLaVA-NeXT (Qwen1.5-72B) 46.25%/45.50%/41.00% 51.75%/52.00%/49.00% 66.75%/67.75%/64.25%
LLaVA-NeXT (Llama3-8B) 44.00%/42.00%/38.75% 46.75%/46.25%/42.75% 53.75%/56.75%/54.75%
IDEFICS-2-8B (Mistral-7B) 44.75%/44.50%/39.75% 46.75%/46.50%/44.00% 58.75%/58.75%/53.75%
DeepSeek-VL (DeepSeek-LLM-7B) 42.25%/40.75%/35.75% 47.00%/47.25%/45.25% 56.75%/58.25%/58.50%
Qwen-VL (Qwen-7B) 44.25%/42.00%/38.75% 48.00%/51.00%/44.75% 61.25%/63.50%/61.25%
3.3 Evaluation on Classification
In this section, we evaluate the performance of LMMs on OBI classification tasks and try to answer: (1) How robust are LMMs in classifying oracle bone characters? Since the intra-class and inter-class similarity of OBIs are sometimes numerically close, we then investigate the capabilities of LMMs in focusing on atypical characteristics and try to answer: (2) How well do LMMs distinguish between different categories of OBIs?
Setup. We sample 500 images across 100 categories from each of the HWOBC, Oracle-50k, and OBI125 datasets as test set. One image is kept as the label in each category and the remaining four as query images. To answer the above questions, we construct three evaluation scenarios: (1) evaluation on oracle bone rubbings and handprinted images, aiming to assess the robustness of LMMs to content qualities. Due to the inadequacy of manual rubbing and the limitation of scanning equipment, there are inevitably various noises in rubbing images, which could greatly affect the classification accuracy; (2) evaluation on different queries (i.e., Yes-or-No question and How question) designed to generate absolute and probability forms of responses, aiming to assess the robustness of LMMs under varied query prompts; (3) evaluation on $\{10,20,30,50,100\}$ classes that encompass mixed image sources and subsets characterized by similar structures, aiming to study the conditions under which LMMs will fail in classification. We thereby reconstruct a variable-quantity mixed test set by sampling an equal number of oracle bone character classes from three datasets. More details are in App. C.3.
Figure 3: Effects of the number of character categories on classification accuracy.
Results. In scenario (1), from Tab. 4, we can observe that the classification results on HWOBC and Oracle-50k datasets are markedly better than those on the OBI125 dataset (surpass by 16.4% and 19.2% for GPT-4o under ‘Yes-or-No’ queries). We believe that this can be attributed to the image quality of OBI in the dataset, which affects the LMM’s perception of character structure. Indeed, the two handprinted OBI datasets, HWOBC and Oracle-50k, are cleaner than the noise-contaminated OBI125 dataset (Sec. C.3). In scenario (2), we find significant performance discrepancies between ‘Yes-or-No’ and ‘How’ queries, which is similar to the other tasks. This highlights the necessity of designing more appropriate textual answer forms and task-specific metrics. Moreover, the Gemini 1.5 series is on par with that of their best competitors (GPT-4v and GPT-4o), which even achieves nearly 90% and 100% accuracy in terms of Acc@1 and Acc@5 metrics respectively. This proves the feasibility of using LMMs for preliminary classification of OBI images. In scenario (3), we select four representative LMMs, i.e., InternVL2-8B, InternVL2-40B, InternVL2-Llama3-76B, and GPT-4v, for evaluation, and obtain three observations from Fig. 3. First, we notice an overall performance improvement compared to the previous experimental setup since the source-mixed test set increases the inter-class differences. Second, the accuracy of these models decreases as the number of classes increases. This shows that current LMMs struggle to perform structural comparisons across a large number of categories, potentially limited by the granularity of their outputs. Third, compared to models with smaller scale, the performance degradation of InternVL2-Llama3-76B and GPT-4v is less pronounced as the number of classification categories increases, indicating that LMMs with larger parameter sizes tend to have better robustness.
Table 5: Performance comparison of LMMs on the OBI retrieval task. Results on OBC306 and OBI-IJDH datasets are delimited by slashes. We report the averaged Recall@1, Recall@3, and Recall@10 on ‘How’ question.
[…]
Gemini 1.5 Pro 0.4316/0.5734 0.225/0.250 0.671/0.688 1.000/1.000 0.644/0.76
Gemini 1.5 Flash 0.3876/0.5344 0.195/0.225 0.635/0.644 0.975/1.000 0.562/0.74
GPT-4v 0.4228/0.5680 0.205/0.225 0.650/0.676 1.000/1.000 0.624/0.74
GPT-4o (ver. 0806) 0.4550/0.6122 0.235/0.250 0.686/0.706 1.000/1.000 0.688/0.80
Qwen-VL-Max (ver. 0809) 0.4223/0.5716 0.190/0.225 0.621/0.638 1.000/1.000 0.664/0.78
GLM-4V 0.3018/0.3867 0.125/0.175 0.495/0.613 0.925/1.000 0.520/0.72
[…]
mPLUG-Owl3 (Qwen2-7B) 0.2814/0.3550 0.075/0.225 0.371/0.588 0.870/0.950 0.384/0.62
MiniCPM-V 2.6 (Qwen2-7B) 0.2863/0.3575 0.070/0.225 0.361/0.581 0.885/0.925 0.368/0.66
InternVL2-Llama3-76B (Llama3-70B) 0.3557/0.4268 0.150/0.250 0.460/0.675 1.000/1.000 0.522/0.72
InternVL2-40B (Nous-Hermes2-Yi-34B) 0.3260/0.3914 0.125/0.250 0.445/0.656 1.000/1.000 0.480/0.68
InternVL2-8B (InternLM2.5-7B) 0.2844/0.3623 0.095/0.225 0.374/0.650 0.925/1.000 0.420/0.68
LLaVA-NeXT (Qwen1.5-72B) 0.3478/0.4143 0.155/0.250 0.468/0.669 1.000/1.000 0.562/0.70
LLaVA-NeXT (Llama3-8B) 0.2793/0.3605 0.075/0.225 0.358/0.606 0.925/1.000 0.348/0.60
IDEFICS-2-8B (Mistral-7B) 0.2844/0.3602 0.090/0.200 0.354/0.619 0.875/1.000 0.402/0.66
DeepSeek-VL (DeepSeek-LLM-7B) 0.2867/0.3450 0.105/0.225 0.343/0.600 0.930/1.000 0.416/0.64
Qwen-VL (Qwen-7B) 0.2883/0.3528 0.080/0.225 0.345/0.588 0.920/0.925 0.422/0.66
3.4 Evaluation on Retrieval
In addition to classification, the retrieval task is also critical for building large-scale OBI image databases. Considering that LMMs combine powerful visual perception and natural language understanding capabilities, they can serve as the demand processing core of OBI retrieval systems. Here, we evaluate the performance of LMMs on OBI retrieval tasks given different types of queries.
Setup. We adopt two query formats (i.e., Yes-or-No and How questions) akin to those employed in the classification tasks. A total of 600 images sampled from OBC306 and OBI-IJDH datasets are used for evaluation. Among each test set, half of the images are used as distractors to simulate a real OBI retrieval scenario. During experiments, OBI images in the same category are retrieved. We take the averaged Recall@k and mean Average Precision (mAP) metrics to quantify the OBI retrieval performance of LMMs in multi-round queries. Note that we use mAP@$|Yes|$ as the evaluation metric in Yes-or-No scenarios, where $|Yes|$ denotes the number of ‘Yes’ answers.
Results. As presented in Tab. 5, most proprietary LMMs can well retrieve relevant OBI contents from test pools except GLM-4v, which is on average 40.48% worse than the other five in terms of Recall@1. Besides, when expanding the search scope to ten candidates, all proprietary LMMs as well as over 63.6% open-source LMMs perfectly retrieve all relevant OBI contents in OBI-IJDH. However, the retrieval performance fell off a cliff in the OBC306 dataset, which we attribute to the same image quality problems encountered with the classification task. As visualized in Fig. 1, such cropped oracle bone rubbings contain various distortion problems, such as stroke-broken, edge-cracked, spindles, and dense white regions, making them harder to tell apart and showing great necessity to add image enhancement measures before these tasks. Furthermore, models with large number of parameters still perform relatively better in this task and How queries receive more precise search results.
Figure 4: Qualitative comparison of deciphering results between two state-of-the-art LMMs, i.e., GPT-4o and Qwen-VL-Max, in a single round of direct questioning. It is noted that neither GPT-4o nor Qwen-VL-Max has fully deciphered these four oracle bone characters.
3.5 Evaluation on Deciphering
Deciphering the oracle bone inscriptions has long been the ultimate goal of this ancient language study. Large multi-modal models have strong cross-modal understanding capabilities, enabling the models to perform multilingual visual-text transformation with linguistic analysis. In this section, we examine the potential OBI deciphering abilities of LMMs by asking: (1) Can LMMs interpret oracle bone characters correctly? (2) How do models behave in the face of different character formation principles (e.g., pictograph, ideogram, and radical-based variants) and character frequencies? (3) What is the reasoning mechanism of LMMs during the OBI deciphering process?
Setup. To evaluate the decipherment performance of LMMs, we construct four evaluation scenarios: (1) deciphering oracle bone characters with different frequency levels $F$. Three test sets (i.e., Tier-1: $F\geq 500$, Tier-2: $100\leq F\leq 500$, Tier-3: $10\leq F\leq 100$) with a total of 40 characters sampled from the HUST-OBS dataset according to the taxonomy in (Liu, 2010; Chen, 2010); (2) deciphering oracle bone characters in pictograph formation. We follow the category definitions in (Chou, 1979; Da WEIH, 2021) and choose 28 representative OBI characters from the EVOBC dataset as test set; (3) deciphering oracle bone characters in ideogram formation. Similar to scenario (2), 22 OBI characters from the EVOBC dataset are selected; (4) deciphering oracle bone characters with structurally similar variants (i.e., radical). We pick out 10 different radicals from the OBI Component 20 dataset (Hu et al., 2024), each accompanied by 5 component-level variants. Fig. 9, Fig. 10, and Fig. 11 in the appendix visualize the sampled test sets for the above scenarios. Different from other approaches (Chang et al., 2022; Guan et al., 2024b) that take the decipherment of OBI as establishing correspondences with modern Chinese characters, we directly use the LMMs to generate a segment of interpretation, as illustrated in Fig. 4. Given the absence of dedicated evaluation criteria for OBI decipherment, we produce text embeddings for all decipherment results and calculate the BERTScore (Zhang et al., 2019) with expert-labelled golden descriptions (i.e., text references). More details are in App. B.3.
Table 6: Results on OBI deciphering task. Given a reference description $x=\left<x_{1},\dots,x_{k}\right>$ and a candidate deciphered result $\hat{x}=\left<\hat{x_{1}},\dots,\hat{x_{k}}\right>$, we use contextual embeddings to represent the tokens, and compute matching using cosine similarity.
Datasets HUST-OBS EVOBC OBI Component 20 Average
[…]
Human (public) 0.4507 0.3884 0.3437 0.4966 0.3627 0.2812 0.3872
[…]
Gemini 1.5 Flash 0.3545 0.3829 0.3567 0.3661 0.3595 0.3648 0.3641
GPT-4v 0.3764 0.3808 0.3596 0.4424 0.3551 0.4060 0.3867
GPT-4o (ver. 0806) 0.3891 0.3510 0.3660 0.4535 0.3893 0.3767 0.3876
Qwen-VL-Max (ver. 0809) 0.2345 0.2273 0.2322 0.2565 0.2533 0.2785 0.2471
GLM-4V 0.2180 0.1638 0.2122 0.2353 0.2451 0.1683 0.2071
[…]
mPLUG-Owl3 (Qwen2-7B) 0.1609 0.1822 0.1897 0.2036 0.1751 0.0684 0.1633
MiniCPM-V 2.6 (Qwen2-7B) 0.1411 0.1694 0.1584 0.1703 0.1529 0.0907 0.1471
Moondream2 (ver. 0728) 0.1129 0.1171 0.1063 0.1297 0.1176 0.1155 0.1165
InternVL2-Llama3-76B (Llama3-70B) 0.2324 0.2139 0.2156 0.2355 0.2123 0.1881 0.2163
InternVL2-40B (Nous-Hermes2-Yi-34B) 0.1966 0.1917 0.1883 0.2007 0.1902 0.1376 0.1842
InternVL2-8B (InternLM2.5-7B) 0.1676 0.1739 0.1340 0.1746 0.1517 0.1670 0.1615
GLM-4V-9B (GLM-4-9B) 0.1291 0.0821 0.0478 0.0835 0.0909 0.0399 0.0789
CogVLM2-Llama3-19B (Llama3-8B) 0.1096 0.0873 0.0852 0.1024 0.1079 0.1192 0.1019
LLaVA-NeXT (Qwen1.5-72B) 0.1501 0.1647 0.1372 0.1468 0.1123 0.0863 0.1329
LLaVA-NeXT (Llama3-8B) 0.0893 0.0873 0.0892 0.0806 0.0813 0.0298 0.0763
IDEFICS-2-8B (Mistral-7B) 0.1639 0.1252 0.0985 0.1601 0.2189 0.1517 0.1531
DeepSeek-VL (DeepSeek-LLM-7B) 0.0458 0.0628 0.0504 0.0324 0.0593 0.0931 0.0573
InternLM-XComposer2-VL (InternLM2-7B) 0.2963 0.3101 0.2660 0.2650 0.1822 0.0626 0.2304
LLaVA-v1.5 (Vicuna-v1.5-13B) 0.1368 0.1669 0.1629 0.1128 0.1214 0.1168 0.1363
LLaVA-v1.5 (Vicuna-v1.5-7B) 0.0677 0.1165 0.0906 0.0924 0.0887 0.1488 0.1008
Qwen-VL (Qwen-7B) 0.1145 0.0912 0.0757 0.1080 0.0635 0.0744 0.0879
Results. For scenario (1), Tab. 6 presents the evaluation results across different character frequencies. We find that the evaluated LMMs achieve on average 0.1905, 0.1898, and 0.1791 on Tier1, Tier2, and Tier3 subsets, respectively, indicating that the deciphering performance on common characters is better than that on rare characters. For scenario (2) and (3), the deciphering results on pictograph are more accurate than ideogram (average BERTScore: 0.2012 > 0.1837), which shows the characteristics of pictographs evolving from drawings and reveals the deciphering mechanism of LMMs, that is, by establishing visual associations between font structure and real objects. For scenario (4), we notice the worst performance (average BERTScore = 0.1636) on the Radical subset compared to other test sets, which means LMMs are impeded in distinguishing and deciphering component-level variants. Furthermore, proprietary LMMs outperform open-source LMMs by an average of 145.9% across all deciphering tasks, suggesting that the open-source LMMs still have a substantial gap to fill for practical application. It is worth noting that GPT-4o, GPT-4v, and Gemini 1.5 Pro have approached or exceeded public-level humans in some scenarios. Surprisingly, InterLM-XComposer2-VL reaches the Top-1 performance among all open-source LMMs (even surpassing InternVL2-Llama3-76B and InternVL2-40B which owns several times of parameters). We suspect that this might be attributed to the specific training datasets (Cai et al., 2024), since the match of data content sometimes outweighs the sheer quantity of parameters (Zhang et al., 2024a). Meanwhile, we observe some highly repetitive outputs (such as ‘The image you’ve provided appears to be a stylized representation of an oracle bone character.’ or other meaningless answers) in the LLaVA-NeXT series and its previous version LLaVA-v1.5, which leads to inferior deciphering results in all scenarios. Fig. 4 provides an example of the deciphering results of GPT-4o and Qwen-VL-Max on four high-frequency oracle bone characters. We also investigate the effect of adding pre-prompt during querying in App. C.4.1. More deciphered results for currently undeciphered OBI are presented in App. C.4.2.
4 Conclusion
We provide comprehensive evaluations of the applicability of LMMs in OBI domain from different perspectives. Overall, there is still considerable room for LMMs to improve when addressing tasks that need domain-specific knowledge. However, taking LMM as a unique technique to assist studies on oracle bone inscriptions is a promising direction, which may significantly reduce learning costs for domain professionals and help them solve OBI-related problems more easily than learning or training traditional deep learning algorithms. Additionally, there are many properties of inputs that affect the performance of OBI processing based on our evaluations, which is worth further exploring. We hope these results can bring insights into the future improvements of LMMs for finer-grained and more robust visual perception. We address the ethical concerns of our research in App. A and discuss the limitations of this work in App. D.
Appendix
Figure 5: Interface of subjective experiments for the recognition task.
Appendix A Ethical Discussions
A.1 Ethical Discussions of Our Research
Our work holds the potential for significant social impact, both positive and negative. Firstly, the main motivation of this work is that the archaeological discoveries of oracle bone inscriptions (OBI) are normally considered as highly serendipitous findings while their interpretation relies heavily on the large volume of manual efforts (from excavation to digital archiving), yet most of their meanings are still unknown, hindering historians and sociologists to further understand the contents of production and life in ancient times. In the early stage of OBI research, the deciphering of oracle bone characters is highly limited, either by referring to ancient books or by working backward from the perspective of font evolution, which was in a fairly stagnant state for much time. Recently, with a large volume of large multi-modal models (LMMs) claiming their powerful cross-modal information processing capability, it is unknown for both computer vision experts and OBI scholars if LMMs can identify and decipher them. In our work, we find while most LMMs are still not completely applicable in some OBI processing tasks that need fine-grained perception and deeper cognition, it exhibits a trend that the performance of the model improves year by year and relies on its design for specific tasks. Our findings offer a promising direction that using LMMs to tackle the OBI problems has a lower knowledge cost than the traditional model-task in one-to-one form on certain occasions. Secondly, our research on the model applicability in OBI tasks provides a necessary understanding of the nature and potential causes of the inferior perception and cognition abilities of LMMs in the knowledge-specific domain. The evaluation of model performance and the subsequent discoveries would spark a broader discussion on the rational use of LMMs to facilitate the study of ancient characters. Our work could serve as a reference point for discussions on developing next-generation LMMs and evaluation metrics for the OBI area. Thirdly, we acknowledge that unregulated and unlimited deciphering results of LMMs may present potential risks in generating misinformation and spreading false cultural values.
A.2 Ethical Discussions of Data Collection
In this section, we detail the ethical concerns that might arise during the experiments. Specifically, we invite five Chinese college students with no background in the ancient Chinese language to represent public-level performance. They are instructed to perform two tasks that are of great interest to OBI scholars, i.e., recognition and deciphering. Note that all participants are clearly informed of the contents in our experiments. Similar to (Chen et al., 2024b; c), we addressed the ethical challenges by obtaining from each subject a signed and informed agreement that they agreed their subjective opinions would be used for non-commercial research, making it equipped with such legal and ethical characteristics.
For the recognition task, we divide the whole subjective experiment into two stages considering the workload and avoiding visual fatigue. In the first stage, participants are asked to describe the provided image contents and determine whether it is an OBI image, thus completing the What and Yes-or-No queries, as depicted in Fig. 5. In the second stage, participants are told to delineate the positions of individual oracle bone characters as they perceive them. We adopt the commonly used toolbox “LabelImg” (Tzutalin, 2015) for annotation. The resulting number of bounding boxes and the corresponding coordinates are used for evaluating the performance under How and Where queries. As for the deciphering task, participants are told to write down their understanding of each oracle bone character, which is then taken to compare with the golden description for alignment evaluation. The user interface is the same as the recognition task (Fig. 5). Overall, it took over a week to complete the whole subjective experiments, where each participant contributed an average of 14.8 hours to attend.
Appendix B More Information on Benchmark Datasets
Figure 6: Sample images of the original oracle bone fragments in the O2BR dataset.
B.1 The Proposed O2BR Dataset
The objective of recognizing oracle bone inscription lies in accurately locating each single character on original oracle bones, rubbings, or fragments, thus assisting downstream classification or deciphering tasks (Xing et al., 2019). In the past decade, a variety of datasets (Guo et al., 2015; Huang et al., 2019; Han et al., 2020; AYNU, 2020; Zhang et al., 2021a; Yue et al., 2022; Wang et al., 2022) have been established for oracle bone character recognition. Among them, Oracle-20k (Guo et al., 2015) was proposed under a controllable lab environment with a total of 20,039 synthesized OBI images in 291 categories. Later, Oracle-50k (Han et al., 2020) was released with more diverse character classes and images. In (Zhang et al., 2021a), a large-scale dataset with character-level annotations, OracleBone-8000, was collected, which contains 7,824 OBI rubbing images scanned from (Guo, 1982) and 128,770 annotated character crops. Similarly, YinQiWenYuan-detection (AYNU, 2020) dataset contributed 9,306 annotated OBI rubbing images with character-level bounding boxes. However, the image content of the above OBI recognition datasets are all handprinted or rubbing form, rather than original oracle bone. More importantly, the majority of datasets either lack the location information for individual oracle characters or do not release the data publicly. To address these problems, we propose the Original Oracle Bone Recognition (O2BR) dataset, which consists of 800 carefully curated original oracle bone images and 4,211 bounding boxes.
B.1.1 Source Content Collection
To curate the O2BR dataset, we began with the initial filtering of available websites using hashtags commonly associated with oracle bone image, such as #oracle bone archaeology, #shell and bone script, and #ancient character. This process allowed us to obtain a set of candidate websites on oracle collection supported by civic organizations, museums, and research institutes around the world. Considering the type, quantity, and usage requirements of the images, we choose to download original oracle bone images from the website [^1] https://openmuseum.tw/objects of open museum led by the Institute of History & Philology, Academia Sinica. Out of the potentially accessible retrieved results, around thirty thousand contained downloadable and visible OBI photos as of July 2024. We utilized a Python-based crawler to download a total of 21,453 raw images. Note that all images downloaded from the website are released under an appropriate Creative Commons (CC) license that allows further editing and redistribution. After that, two graduate students with oracle research backgrounds were involved in removing small fragments that do not contain complete oracle bone characters. Finally, 800 original oracle bone images are selected as the annotation source. Fig. 6 presents several samples of this dataset. Tab. 7 provides detailed statistics.
Table 7: Statistics of the proposed O2BR dataset.
Attribute Value (in pixels)
Minimum Width 167
Maximum Width 2167
Minimum Height 168
Maximum Height 2664
Average Width 1529
Average Height 1192
Avg. number of bounding boxes/image $\approx$ 5.264
B.1.2 Annotation Process
Some studies (Li et al., 2020; Diao et al., 2023) suggested the use of automatic object detection methods to save manual annotation effort. However, considering the particularity of the image content in this dataset and the lack of relevant trainable datasets, we adopted the commonly used toolbox “LabelImg” (Tzutalin, 2015) to conduct the character-level annotation. Fig. 7(a) shows the annotation interface. Two OBI domain experts participated in this process. Besides, to avoid duplication of evaluation tasks in OBI-Bench, we merely required the annotator to frame the location of single oracle bone character. Afterwards, a senior OBI domain experts was invited to proofread the annotations. If there is a disagreement between two experts, the senior expert will re-annotate the image and determine the final annotation, which helps to guarantee the quality of our proposed dataset.
Figure 7: The illustration of the annotation interfaces for the O2BR dataset (coordinate) on recognition task, and the OBI-rejoin dataset (adjacency) on rejoining task.
B.2 The Proposed OBI-rejoin Dataset
Table 8: Statistics of the proposed OBI-rejoin dataset.
Attribute Value (in pixels)
map fragments
Minimum Width 64 46
Maximum Width 1268 909
Minimum Height 107 29
Maximum Height 2913 1273
Average Width 370 265
Average Height 465 313
Avg. number of fragments/map = 2.415
Figure 8: Sample images of the original OBI rejoining results (left) as well as the separated fragments (right) in the OBI-rejoin dataset.
In the early days, the most important reference in OBI domain is the “Collections of oracle bone inscriptions” book (Guo, 1982), which contains nearly 2,000 groups of OBI rejoining results. Other representative books such as “Catalog of Oracle Bone Rejoinings"  (Tsai, 1999), “Zui Gu: Research on Oracle Bone Rejoinings" (Lin, 2008), “The Fifth Collection of Oracle Bone Rejoinings" (Huang, 2019), and “Compilation of Oracle Bone Inscriptions: Reconstruction and Interpretation" (Zhang, 2020) involve the achievements of over 4,000 groups of OBI rejoining results in the past 30 years. However, there is no publicly available rejoining dataset for any type of OBIs in digital image form. Recently, researchers collected a real-world dataset for rejoining Oracle Bone fragments (Zhang et al., 2022), which consists of 998 oracle bone rubbing images with manually outlined top and bottom marginal areas. Unfortunately, its proprietary nature limits its use by the community. This motivates us to build the OBI-rejoin, an OBI dataset comprising 200 complete OBI pieces with 483 rejoinable fragments, shown in Fig. 8. Tab. 8 lists the statistical information of OBI-rejoin dataset.
B.2.1 Source Content Collection
We used a Python-based crawler to scrape metadata from the official website [^2] https://www.xianqin.org/blog/archives/category/jgw_study/jgw_zhuihe of the Pre-Qin Research Office at the Institute of History, Chinese Academy of Social Sciences. This process covered 217 pages and encompassed the time frame from December 2005 to July 2024, resulting in 4,249 OBI images [^3] We extend our deepest gratitude to the frontline researchers and scholars involved in the meticulous collation and proofreading of the oracle bone inscriptions.. To ensure data quality, we conducted an expert examination that manually excluded hand-drawn contour images as well as extraneous contents, retaining only inked rubbings and original oracle bone forms as the image sources of the OBI-rejoin dataset. As a result, a total of 200 OBI images covering two forms of OBI (i.e., original and inked) with different character styles and resolutions are selected.
B.2.2 Annotation Process
Here, we refer to the OBI rejoining results as “map” to reflect its special role in the construction process of OBI-rejoin dataset. Three domain experts in OBI research and one senior OBI scholar were involved in the annotation process of OBI-rejoin. Each annotator was equipped with a PC with commercial photo editing software. Fig. 7(b) illustrates the annotation interface. To precisely reproduce the authentic conditions of oracle bone fractures, the annotator manually tears “maps” on appropriate borders utilizing their domain expertise to induce central or peripheral fractures. A double-check procedure was applied after the annotation to ensure quality and rationality. Overall, it took each annotator approximately 3 hours to finish annotation, and the whole construction process took over a week to complete. A lot of effort and discussion were required in order to manually retrieve suitable rejoining results from the crawled metadata. To the best of our knowledge, OBI-rejoin is the first open-source oracle bone inscription rejoining dataset. We hope that OBI-rejoin can facilitate future development of advancing OBI rejoining algorithms and serve as a benchmark dataset to evaluate the performance of rejoining methods.
Figure 9: Visualization of 40 oracle bone characters with different frequency levels.
Figure 10: Visualization of 28 and 22 oracle bone characters in pictograph and ideogram formation, respectively.
Figure 11: Visualization of 10 different radicals with their structural variants as well as the corresponding modern Chinese characters.
B.3 Collecting Golden Descriptions for Deciphering Task
The existing OBI deciphering works merely decipher oracle bone characters by linking them to modern Chinese characters (Chang et al., 2022; Guan et al., 2024b), which does not constitute a true decipherment. With the historical and cultural changes, many oracle bone characters can no longer find correspondences from modern Chinese characters. Thus, we turn to the most traditional dictionary form and extract golden descriptions from an authoritative Chinese character database (CUHK, 2014).
Appendix C Additional Experimental Results
C.1 Benchmark Candidates
Table 9: A brief view of 17 open-source LMMs evaluated in the OBI-Bench in chronological order.
[…]
${}^{\textit{24.08}}$mPLUG-Owl3-7B (Ye et al., 2024) SigLIP-400M 384 MLP projector Qwen2-7B decoder-only
${}^{\textit{24.08}}$iniCPM-V 2.6-8B (Yao et al., 2024) SigLIP-400M 224 Perceiver resampler Qwen2-7B decoder-only
${}^{\textit{24.08}}$oondream2-1.6B (Moondream.ai, 2024) SigLIP-400M 378 MLP projector Phi-1.5 decoder-only
${}^{\textit{24.07}}$nternVL2-Llama3-76B (Chen et al., 2024a) InternViT-6B 448 MLP projector Hermes-2-Theta-Llama-3-70B decoder-only
${}^{\textit{24.07}}$nternVL2-40B (Chen et al., 2024a) InternViT-6B 448 MLP projector Nous-Hermes-2-Yi-34B decoder-only
${}^{\textit{24.07}}$nternVL2-8B (Chen et al., 2024a) InternViT-300M 448 MLP projector InternLM2.5-7B decoder-only
${}^{\textit{24.06}}$-4V-9B (GLM et al., 2024) EVA2-CLIP-E/14+ 224 MLP projector GLM-4-9B decoder-only
${}^{\textit{24.05}}$ogVLM2-LLaMA3-Chat-19B (Wang et al., 2023) EVA2-CLIP-E 224 MLP projector Llama-3-8B-Instruct decoder-only
${}^{\textit{24.05}}$aVA-NeXT-72B (Li et al., 2024a) CLIP-ViT-L/14 336 MLP projector Qwen1.5-72B decoder-only
${}^{\textit{24.05}}$aVA-NeXT-8B (Li et al., 2024a) CLIP-ViT-L/14 336 MLP projector Llama-3-8B-Instruct decoder-only
${}^{\textit{24.05}}$-2-8B Laurençon et al. (2024) SigLip-400M 384 MLP projector Mistral-7B decoder-only
${}^{\textit{24.03}}$eepSeek-VL-7B (Lu et al., 2024) SAM-B&SigLIP-L 1024&384 MLP projector DeepSeek-LLM-7B decoder-only
${}^{\textit{24.01}}$nternLM-XComposer2-VL-7B (Dong et al., 2024) CLIP-ViT-L/14 336 Partial LoRA InternLM2-7B decoder-only
${}^{\textit{23.10}}$aVA-v1.5-13B (Liu et al., 2024a) CLIP-ViT-L/14 336 MLP projector Vicuna-v1.5-13B decoder-only
${}^{\textit{23.10}}$aVA-v1.5-7B (Liu et al., 2024a) CLIP-ViT-L/14 336 MLP projector Vicuna-v1.5-7B decoder-only
${}^{\textit{23.08}}$wen-VL (Bai et al., 2023) CLIP-ViT-G/14 448 Cross-attention Qwen-7B decoder-only
To ensure the comprehensiveness and timeliness of the results, we select 23 up to date and prevailing LMMs for evaluation. The proprietary LMMs include Gemini 1.5 Pro (Reid et al., 2024), Gemini 1.5 Flash (Reid et al., 2024), GPT-4v (Achiam et al., 2023), GPT-4o (OpenAI, 2024a), Qwen-VL-Max (Bai et al., 2023), and GLM-4v (GLM et al., 2024). The open-source LMMs include xGen-MM-instruct-interleave (Xue et al., 2024), mPLUG-Owl3-7B (Ye et al., 2024), MiniCPM-V 2.6-8B (Yao et al., 2024), Moondream2-1.6B (Moondream.ai, 2024), InternVL2-Llama3-76B (Chen et al., 2024a), InternVL2-40B (Chen et al., 2024a), InternVL2-8B (Chen et al., 2024a), GLM-4V-9B (GLM et al., 2024), CogVLM2-LLaMA3-Chat-19B (Wang et al., 2023), LLaVA-NeXT-72B (Li et al., 2024a), LLaVA-NeXT-8B (Li et al., 2024a), IDEFICS-2-8B Laurençon et al. (2024), DeepSeek-VL-7B (Lu et al., 2024), InternLM-XComposer2-VL-7B (Dong et al., 2024), LLaVA-v1.5-13B (Liu et al., 2024a), LLaVA-v1.5-7B (Liu et al., 2024a), and Qwen-VL (Bai et al., 2023). Tab. 9 summarizes and compares the vision and language architectures of open-source LMMs.
Additionally, considering the cultural uniqueness of OBI, and to ensure the representativeness of human performance on the OBI-Bench, we recruited five Chinese college students with no background in ancient Chinese language to reflect public cognition. Initially, we instructed all participants to have a clear and consistent understanding of all question-answer patterns by familiarizing themselves with the tasks through exposure to similar cases. Note that we conduct user studies in the same order of tasks as we test LMMs to avoid interference between tasks. Given the inherent variability of LMMs, identical prompts can yield responses. To address the impact of such situations on our evaluation, we implemented a 5-round average strategy.
C.2 Additional Details of Evaluation on Recognition
C.2.1 Evaluation Criteria
For What question, we calculate the cosine similarity between the embeddings of output answers and the golden descriptions. Specifically, we use the Sentence-Transformers python library and the all-MiniLM-L6-v2444https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 model, which maps sentences to a 384 dimensional dense vector space.
For How question, we use the relative counting error metric to measure the performance of different LMMs:
$\text{MRE}=\frac{1}{N}\sum_{i=1}^{N}\frac{\left|C_{i}^{truth}-C_{i}^{pred}\right|}{C_{i}^{truth}}$ (1)
where $N$ is the number of evaluated OBI images. $C_{i}^{truth}$ and $C_{i}^{pred}$ denote the actual number of oracle bone characters in the image and the predicted number of oracle bone characters respectively.
Table 10: Impact of the judgment bias. Answer accuracy of LMMs on Yes-or-No questions.
LMM Yes-or-No
Gemini 1.5 Pro 96.15%
Gemini 1.5 Flash 93.37%
GPT-4v 95.13%
GPT-4o 98.68%
Qwen-VL-Max 92.26%
GLM-4V 86.68%
[…]
mPLUG-Owl3 (Qwen2-7B) 87.68%
MiniCPM-V 2.6 (Qwen2-7B) 82.21%
InternVL2-Llama3-76B (Llama3-70B) 93.59%
InternVL2-40B (Nous-Hermes2-Yi-34B) 92.24%
InternVL2-8B (InternLM2.5-7B) 86.55%
GLM-4V-9B (GLM-4-9B) 81.62%
CogVLM2-Llama3-19B (Llama3-8B) 88.68%
LLaVA-NeXT (Qwen1.5-72B) 94.77%
LLaVA-NeXT (Llama3-8B) 85.50%
IDEFICS-2-8B (Mistral-7B) 87.55%
DeepSeek-VL (DeepSeek-LLM-7B) 84.37%
InternLM-XComposer2-VL (InternLM2-7B) 86.88%
LLaVA-v1.5 (Vicuna-v1.5-13B) 79.93%
LLaVA-v1.5 (Vicuna-v1.5-7B) 75.58%
Qwen-VL (Qwen-7B) 84.57%
C.2.2 Extended Results on Yes-or-No Question
Considering that the test set used in the recognition task is composed solely of authentic oracle bone rubbings or fragments, it may be affected by the response preferences of LMMs in Yes-or-No questions. In this section, we take a deeper analysis of the Yes-or-No judgment ability of LMMs, that whether these models can get similar accuracy on questions that should be answered with Yes, as those should be replied as No. Specifically, we re-curate a content-mixed test set, which consists of 1,000 OBI images and other images containing ancient texts (such as the inscription on the bronzes, characters engraved on bamboo slips, and characters written on silk). As a result, the proportion of OBI images and non-OBI images is 1:1. As reported in Tab. 10, the performance of GLM-4V and GPM-4V-9B improves greatly compared to the results in the O2BR and YinQiWenYuandetection datasets, while the other LMMs perform at a similar level.
C.3 Additional Details of Evaluation on Classification
To explore the effect of OBI image quality on classification accuracy, we further select two classical image quality assessment (IQA) metrics, namely NIMA (Talebi & Milanfar, 2018) and CNNIQA (Kang et al., 2014), to measure the average image quality within each test sets. As listed in Tab. 11, the numerical results of both metrics show that the image quality of HWOBC and Oracle-50k datasets is better than that of OBI125. A similar conclusion to the main paper can be drawn that enhancing the image quality or applying denoising operation can improve the OBI classification accuracy of LMMs.
Table 11: Image quality comparison between two handprinted OBI datasets and one cropped from real rubbings.
Metric HWOBC Oracle-50k OBI125
[…]
Table 12: Effect of pre-prompt on the OBI deciphering task. We select Gemini 1.5 Pro, GPT4o, Qwen-VL-Max, and GLM-4V for evaluation.
Datasets HUST-OBS EVOBC OBI Component 20 Average
[…]
Gemini 1.5 Pro 0.3766 0.3834 0.3589 0.4226 0.3696 0.3750 0.3810
Gemini 1.5 Pro+Role 0.3824 0.3863 0.3603 0.4347 0.3819 0.3738 0.3866
Gemini 1.5 Pro+Role+Case 0.3837 0.3867 0.3626 0.4386 0.3779 0.3722 0.3870
[…]
GPT-4o+Role 0.3925 0.3525 0.3557 0.4585 0.3817 0.3729 0.3856
GPT-4o+Role+Case 0.3922 0.3586 0.3688 0.4545 0.3829 0.3783 0.3892
[…]
Qwen-VL-Max+Role 0.2447 0.2535 0.2533 0.2585 0.2401 0.2796 0.2549
Qwen-VL-Max+Role+Case 0.2518 0.2778 0.2617 0.2830 0.2457 0.2883 0.2681
[…]
GLM-4V+Role 0.1977 0.1984 0.1989 0.2554 0.2448 0.1689 0.2107
GLM-4V+Role+Case 0.2187 0.1968 0.2137 0.2585 0.2487 0.1710 0.2179
C.4 Additional Details of Evaluation on Deciphering
C.4.1 Effect of Pre-Prompt
Existing studies (Chen et al., 2023; Zhou et al., 2023) indicate that providing specific instructions, contextual information, or role settings allows the model to understand the user’s needs and expectations and thus provide a more accurate and relevant response. We thereby employ two schemes (i.e., role assignment and case instruction) to investigate the effect of pre-prompt.
Figure 12: Example of three oracle bone character images that are taken as case instructions along with their golden descriptions.
Prompt Template for Role Assignment:
#System: You are a senior oracle bone researcher who excels in interpreting ancient texts. Please help me to predict the meaning of the undeciphered oracle bone characters reasonably and accurately.
#User: What is the meaning of this oracle bone character? <image>
Prompt Template for Case Instruction: We further randomly select three cases (OBI images with correct descriptions) as case instructions (Fig. 12) after role assignment.
#System: Here are three examples of deciphered oracle bone characters.
#System: (1) <image1> The oracle bone character in this image means the blood, which also represents the blood of animals offered to the gods in sacrifices.
#System: (2) <image2> The oracle bone character in this image refers to the pattern and texture of the bamboo tubes merged together on a musical instrument.
#System: (3) <image3> The oracle bone character in this image means the armpit, with the small dots indicating the armpits. It is like a person sweating in his armpits.
#System: Please refer to the case form given above for the following answers.
#User: What is the meaning of this oracle bone character? <image>
In Tab. 12, we compare the deciphering performance of LMMs with or without pre-prompt operations. It can be observed that adding role assignment and case instruction slightly improves the accuracy and reasonableness of decipherment results for Gemini 1.5 Pro, GPT-4o, Qwen-VL-Max, and GLM-4V by an average of 2.02%, 0.40%, 8.50%, and 5.21% respectively. This demonstrates the effectiveness and necessity of prompt engineering in refining the answers of LMMs and avoiding meaningless outputs.
C.4.2 Deciphering Results on Unknown OBIs
We provide the deciphering results of LMMs for previously undeciphered five oracle bone characters from the HUST-OBS dataset (Wang et al., 2024a), whose interpretation is central to the core of Pre-Qin cultural and historical research, in Fig. 14, Fig. 15, Fig. 16, Fig. 17, and Fig. 18. Different from existing decipherment studies (Zhang et al., 2021b; Guan et al., 2024b) that decipher oracle bone characters as characters in subsequent writing systems or pseudo-modern Chinese characters, we directly leverage the multi-modal capabilities of LMMs to decipher them as a description in English. Since the correctness of an oracle decipherment result is usually determined by peer review, we plan to make the code used in OBI-Bench and a comprehensive collection encompassing more decipherment outcomes publicly available. We hope this contribution will assist OBI researchers in advancing the study of ancient languages.
Table 13: Taxonomy of LMMs categorized by the affiliation of the first author.
Model Affiliation of the First Author Country
Proprietary LMMs:
[…]
Gemini 1.5 Flash Google Inc. USA
GPT-4v OpenAI Inc. USA
GPT-4o OpenAI Inc. USA
Qwen-VL-Max Alibaba Group China
GLM-4V Zhipu AI Inc. China
Open-source LMMs:
[…]
mPLUG-Owl3 (Qwen2-7B) Alibaba Group China
MiniCPM-V 2.6 (Qwen2-7B) ModelBest Inc. China
Moondream2 (ver. 0728) M87 Labs USA
InternVL2-Llama3-76B (Llama3-70B) Shanghai AI Lab China
InternVL2-40B (Nous-Hermes2-Yi-34B) Shanghai AI Lab China
InternVL2-8B (InternLM2.5-7B) Shanghai AI Lab China
GLM-4V-9B (GLM-4-9B) Zhipu AI Inc. China
CogVLM2-Llama3-19B (Llama3-8B) Tsinghua University China
LLaVA-NeXT (Qwen1.5-72B) University of Wisconsin-Madison USA
LLaVA-NeXT (Llama3-8B) University of Wisconsin-Madison USA
IDEFICS-2-8B (Mistral-7B) Hugging Face Inc. USA
DeepSeek-VL (DeepSeek-LLM-7B) DeepSeek-AI Inc. China
InternLM-XComposer2-VL (InternLM2-7B) Shanghai AI Laboratory China
LLaVA-v1.5 (Vicuna-v1.5-13B) University of Wisconsin–Madison USA
LLaVA-v1.5 (Vicuna-v1.5-7B) University of Wisconsin–Madison USA
Qwen-VL (Qwen-7B) Alibaba China
Figure 13: Comparison of the average ranking of LMMs under region-based taxonomy.
C.5 Impact of Potential Culture Bias
We further investigate the impact of different cultural backgrounds on OBI processing performance. Tab. 13 classifies LMMs by the affiliation of the first author. Among all evaluated LMMs, there are 12 LMMs from China and 11 LMMs from the USA. To ensure the fairness of the results, we exclude the earliest LMM (Qwen-VL) to make the number of models the same. Given the differences in evaluation criteria and scales of different tasks, we report the average ranking of LMMs according to the region-based taxonomy, shown in Fig. 13. We find that cultural differences did not particularly affect the overall performance of the models. Besides, the LMMs originating from China still have some distance behind those from the United States.
Appendix D Limitations
Although our study provides a comprehensive evaluation of LMMs in OBI tasks, there are several potential limitations acknowledged below. First, the images in our test sets are limited in scope with regard to size, content types for each task, and resolutions, which constrains the applicability of the OBI-Bench. We merely sample partial images from YinQiWenYuandetection, HWOBC, Oracle-50k, OBI125, OBC306, OBI-IJDH, EVOBC, OBI Component 20, and HUST-OBS datasets as test sets. Our self-collected O2BR and OBI-rejoin datasets contain 800 and 483 images. Besides, in the deciphering task, we selected only 140 oracle bone characters for evaluation, which is considerably fewer than the number of deciphered oracle bone characters currently known. Another limitation of our work is that the query-answer forms used in our benchmark are relatively simple. There is no extra information that gives the LMMs clues to the intent of the image. The lack of such context hinders performance in cases such as OBI deciphering where context (role assignment and case instruction), has been shown to aid in image interpretation (App. C.4.1). Furthermore, coarse-grained perception perspectives such as What query in the recognition task and deciphering task involve subjectivity and should be based on general consensus while defining golden descriptions.
These limitations highlight the need for related future research. We encourage the community to view our work as a starting point and extend the evaluations and analysis to further uncover potential capabilities of LMMs not only in the OBI domain but also in other ancient character processing domains and design possible task-oriented practical strategies accordingly.
Appendix E Future Possibilities
Apart from the comments for possible next steps of research related to LMMs and OBI that have already been given, this section is devoted to the extension for some of them and then more topics with good potential based upon our understanding and rethinking for the field.
New possibilities of LMM in aiding OBI processing advancement can be tried in a number of major directions:
  • Focus on Robust Preprocessing Pipelines: Developing preprocessing techniques that can enhance low-quality or noisy OBI data, such as denoising, fragment reconstruction, or super-resolution methods, may significantly improve LMM’s performance on these tasks.
  • Domain-Specific Fine-Tuning with Interdisciplinary Collaboration: Fine-tuning models on domain-specific datasets can help bridge the gap in domain knowledge. Integrating visual and textual modalities can enrich the understanding of oracle bone inscription, thus enhancing the overall interpretative capability. Bridging the gap between computer science and humanities disciplines is crucial for effective application of LMMs in OBI studies. Collaborating with historians and linguists to create multi-modal annotated OBI datasets may benefit this effort.
  • Interactive Question Answering Systems: Future applications could include interactive question-answering systems where users can inquire about oracle bone inscription meanings and historical contexts in natural language, leveraging large multi-modal models for intuitive and accessible interactions.
It is hoped that these can provide actionable ideas for researchers and would trigger further discussion, and more importantly, new exploration in this field.
Figure 14: A qualitative comparison for the deciphered results of 7 representative LMMs on a genuine undeciphered oracle bone character (I).
Figure 15: A qualitative comparison for the deciphered results of 7 representative LMMs on a genuine undeciphered oracle bone character (II).
Figure 16: A qualitative comparison for the deciphered results of 7 representative LMMs on a genuine undeciphered oracle bone character (III).
Figure 17: A qualitative comparison for the deciphered results of 7 representative LMMs on a genuine undeciphered oracle bone character (IV).
Figure 18: A qualitative comparison for the deciphered results of 7 representative LMMs on a genuine undeciphered oracle bone character (V).

### v2

@@
 OBI-Bench includes 5,523 meticulously collected diverse-sourced images, covering five key domain problems: recognition, rejoining, classification, retrieval, and deciphering.
-These images span centuries of archaeological findings and years of research by front-line scholars, comprising multi-stage font appearances from excavation to synthesis, such as original oracle bone, inked rubbings, oracle bone fragments, cropped single character, and handprinted character.
+These images span centuries of archaeological findings and years of research by front-line scholars, comprising multi-stage font appearances from excavation to synthesis, such as original oracle bone, inked rubbings, oracle bone fragments, cropped single characters, and handprinted characters.
 Unlike existing benchmarks, OBI-Bench focuses on advanced visual perception and reasoning with OBI-specific knowledge, challenging LMMs to perform tasks akin to those faced by experts.
@@
 Even the latest versions of GPT-4o, Gemini 1.5 Pro, and Qwen-VL-Max are still far from public-level humans in some fine-grained perception tasks.
-However, they perform at a level comparable to untrained humans in deciphering task, indicating remarkable capabilities in offering new interpretative perspectives and generating creative guesses.
+However, they perform at a level comparable to untrained humans in deciphering tasks, indicating remarkable capabilities in offering new interpretative perspectives and generating creative guesses.
 We hope OBI-Bench can facilitate the community to develop domain-specific multi-modal foundation models towards ancient language research and delve deeper to discover and enhance these untapped potentials of LMMs.
@@
 Nowadays, the emergent large multi-modal models (LMMs) have brought opportunities for solving multidisciplinary tasks with powerful visual perception, understanding, and reasoning abilities (Yue et al., 2024; Caffagni et al., 2024; OpenAI, 2024b).
-Meanwhile, the ability of LMMs remains unclear on fine-grained perception and cognition (i.e., high-level perception), which play significant roles in interpreting ancient texts and its associated tasks on image material pre-processing.
+Meanwhile, the ability of LMMs remains unclear on fine-grained perception and cognition (i.e., high-level perception), which play significant roles in interpreting ancient texts and their associated tasks on image material pre-processing.
 The perception at different granularities is strongly associated with a wide range of OBI tasks, such as distinguishing noise from real oracle bone character recognition (Wang et al., 2022), marginal sealability checking in rejoining (Zhang et al., 2021c) and component-level structural decomposition in retrieval (Hu et al., 2024).
@@
 We showcase some empirical findings here: 1) LMMs still have much room for improvement in fine-grained OBI recognition tasks such as quantity detection or locating; 2) LMMs are not sensitive enough to local information such as the borders of the fragmented margin to directly meet the requirements of rejoining.
-Meanwhile, part of LMMs can help OBI scholars narrow down the range of qualified fragments (GPT-4o and Gemini 1.5 Pro achieves an average performance of 77.57% on the Acc@10 metric); 3) LMMs can be applied in classification and retrieval tasks, especially for cleaner handprinted datasets; 4) partial LLMs rival the performance of public-level humans in deciphering tasks and exhibit remarkable visual reasoning ability in interpreting those genuine undeciphered oracle bone characters.
-The full list of our findings from different OBI tasks is in the Sec.
+Meanwhile, part of LMMs can help OBI scholars narrow down the range of qualified fragments (GPT-4o and Gemini 1.5 Pro achieve an average performance of 77.57% on the Acc@10 metric); 3) LMMs can be applied in classification and retrieval tasks, especially for cleaner handprinted datasets; 4) partial LMMs rival the performance of public-level humans in deciphering tasks and exhibit remarkable visual reasoning ability in interpreting those genuine undeciphered oracle bone characters.
+The full list of our findings from different OBI tasks is in Sec.
 3.
@@
 Covering Multi-stage Font Appearances.
-To cover diverse appearances of the oracle bones since their excavation, we collect multi-sourced images for each task, as depicted in Tab.
+To cover the diverse appearances of the oracle bones since their excavation, we collect multi-sourced images for each task, as depicted in Tab.
 1.
@@
 We use mean relative error (MRE) as the metric.
-(4) evaluating via the Where query to achieve finer-grained perception (i.e., dense oracle bone character locating), designed to more effectively uncover the practicability of LMMs in real OBI processing pipeline.
+(4) evaluating via the Where query to achieve finer-grained perception (i.e., dense oracle bone character locating), designed to more effectively uncover the practicability of LMMs in real OBI processing pipelines.
 Mean intersection over union (mIoU) is used as the metric.
@@
 We speculate it is due to the answer preference bias (Wu et al., 2024).
-Deeper analysis is in App.
+A deeper analysis is in App.
 C.2.2.
@@
 2 indicate that GLM-4V and InternVL2-Llama3-76B exhibit the best performance within their respective division groups.
-Additionally, we observe some abnormally large values (>50) or repeated values from Gemini 1.5 Flash, xGen-MM-4B, LLaVA-NeXT-72B, LLaVA-v1.5-13B, and Qwen-VL that results in their relatively large MRE (Zhou et al., 2024).
+Additionally, we observe some abnormally large values (>50) or repeated values from Gemini 1.5 Flash, xGen-MM-4B, LLaVA-NeXT-72B, LLaVA-v1.5-13B, and Qwen-VL that result in their relatively large MRE (Zhou et al., 2024).
 In scenario (4), all evaluated LMMs consistently underperform.
@@
 Setup.
-Given the absence of publicly available rejoining datasets in OBI domain, we create a dataset consisting of 200 complete OBI pieces across two appearances (i.e., original and inked oracle bone fragments) with 483 rejoinable fragments.
+Given the absence of publicly available rejoining datasets in the OBI domain, we create a dataset consisting of 200 complete OBI pieces across two appearances (i.e., original and inked oracle bone fragments) with 483 rejoinable fragments.
 More details are in App.
@@
 This shows that current LMMs struggle to perform structural comparisons across a large number of categories, potentially limited by the granularity of their outputs.
-Third, compared to models with smaller scale, the performance degradation of InternVL2-Llama3-76B and GPT-4v is less pronounced as the number of classification categories increases, indicating that LMMs with larger parameter sizes tend to have better robustness.
+Third, compared to models with smaller scales, the performance degradation of InternVL2-Llama3-76B and GPT-4v is less pronounced as the number of classification categories increases, indicating that LMMs with larger parameter sizes tend to have better robustness.
 Table 5: Performance comparison of LMMs on the OBI retrieval task.
@@
 1, such cropped oracle bone rubbings contain various distortion problems, such as stroke-broken, edge-cracked, spindles, and dense white regions, making them harder to tell apart and showing great necessity to add image enhancement measures before these tasks.
-Furthermore, models with large number of parameters still perform relatively better in this task and How queries receive more precise search results.
+Furthermore, models with a large number of parameters still perform relatively better in this task and How queries receive more precise search results.
 Figure 4: Qualitative comparison of deciphering results between two state-of-the-art LMMs, i.e., GPT-4o and Qwen-VL-Max, in a single round of direct questioning.
@@
 Surprisingly, InterLM-XComposer2-VL reaches the Top-1 performance among all open-source LMMs (even surpassing InternVL2-Llama3-76B and InternVL2-40B which owns several times of parameters).
-We suspect that this might be attributed to the specific training datasets (Cai et al., 2024), since the match of data content sometimes outweighs the sheer quantity of parameters (Zhang et al., 2024a).
+We suspect that this might be attributed to the specific training datasets (Cai et al., 2024) since the match of data content sometimes outweighs the sheer quantity of parameters (Zhang et al., 2024a).
 Meanwhile, we observe some highly repetitive outputs (such as ‘The image you’ve provided appears to be a stylized representation of an oracle bone character.’ or other meaningless answers) in the LLaVA-NeXT series and its previous version LLaVA-v1.5, which leads to inferior deciphering results in all scenarios.
@@
 D.
+Acknowledgments
+This work was supported in part by the National Key R&D Program of China under Grant 2021YFE0206700, in part by the National Social Science Fund of China under Grant 24Z300404220, and in part by the Shanghai Office of Philosophy and Social Science under Grant 2023BYY003.
 Appendix
@@
 Note that all participants are clearly informed of the contents in our experiments.
-Similar to (Chen et al., 2024b; c), we addressed the ethical challenges by obtaining from each subject a signed and informed agreement that they agreed their subjective opinions would be used for non-commercial research, making it equipped with such legal and ethical characteristics.
+Similar to (Chen et al., 2024b; c; d), we addressed the ethical challenges by obtaining from each subject a signed and informed agreement that they agreed their subjective opinions would be used for non-commercial research, making it equipped with such legal and ethical characteristics.
 For the recognition task, we divide the whole subjective experiment into two stages considering the workload and avoiding visual fatigue.
@@
 5).
-Overall, it took over a week to complete the whole subjective experiments, where each participant contributed an average of 14.8 hours to attend.
+Overall, it took over a week to complete the whole subjective experiment, where each participant contributed an average of 14.8 hours to attend.
 Appendix B More Information on Benchmark Datasets
@@
 Similarly, YinQiWenYuan-detection (AYNU, 2020) dataset contributed 9,306 annotated OBI rubbing images with character-level bounding boxes.
-However, the image content of the above OBI recognition datasets are all handprinted or rubbing form, rather than original oracle bone.
+However, the image contents of the above OBI recognition datasets are all handprinted or rubbing form, rather than original oracle bone.
 More importantly, the majority of datasets either lack the location information for individual oracle characters or do not release the data publicly.
@@
 Two OBI domain experts participated in this process.
-Besides, to avoid duplication of evaluation tasks in OBI-Bench, we merely required the annotator to frame the location of single oracle bone character.
-Afterwards, a senior OBI domain experts was invited to proofread the annotations.
+Besides, to avoid duplication of evaluation tasks in OBI-Bench, we merely required the annotator to frame the location of a single oracle bone character.
+Afterward, a senior OBI domain expert was invited to proofread the annotations.
 If there is a disagreement between two experts, the senior expert will re-annotate the image and determine the final annotation, which helps to guarantee the quality of our proposed dataset.
@@
 The existing OBI deciphering works merely decipher oracle bone characters by linking them to modern Chinese characters (Chang et al., 2022; Guan et al., 2024b), which does not constitute a true decipherment.
-With the historical and cultural changes, many oracle bone characters can no longer find correspondences from modern Chinese characters.
+With the historical and cultural changes, many oracle bone characters can no longer find correspondence with modern Chinese characters.
 Thus, we turn to the most traditional dictionary form and extract golden descriptions from an authoritative Chinese character database (CUHK, 2014).
@@
 ${}^{\textit{23.08}}$wen-VL (Bai et al., 2023) CLIP-ViT-G/14 448 Cross-attention Qwen-7B decoder-only
-To ensure the comprehensiveness and timeliness of the results, we select 23 up to date and prevailing LMMs for evaluation.
+To ensure the comprehensiveness and timeliness of the results, we select 23 up-to-date and prevailing LMMs for evaluation.
 The proprietary LMMs include Gemini 1.5 Pro (Reid et al., 2024), Gemini 1.5 Flash (Reid et al., 2024), GPT-4v (Achiam et al., 2023), GPT-4o (OpenAI, 2024a), Qwen-VL-Max (Bai et al., 2023), and GLM-4v (GLM et al., 2024).
@@
 9 summarizes and compares the vision and language architectures of open-source LMMs.
-Additionally, considering the cultural uniqueness of OBI, and to ensure the representativeness of human performance on the OBI-Bench, we recruited five Chinese college students with no background in ancient Chinese language to reflect public cognition.
+Additionally, considering the cultural uniqueness of OBI, and to ensure the representativeness of human performance on the OBI-Bench, we recruited five Chinese college students with no background in the ancient Chinese language to reflect public cognition.
 Initially, we instructed all participants to have a clear and consistent understanding of all question-answer patterns by familiarizing themselves with the tasks through exposure to similar cases.
@@
 For What question, we calculate the cosine similarity between the embeddings of output answers and the golden descriptions.
-Specifically, we use the Sentence-Transformers python library and the all-MiniLM-L6-v2444https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 model, which maps sentences to a 384 dimensional dense vector space.
+Specifically, we use the Sentence-Transformers python library and the all-MiniLM-L6-v2444https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 model, which maps sentences to a 384-dimensional dense vector space.
 For How question, we use the relative counting error metric to measure the performance of different LMMs:
@@
 Considering that the test set used in the recognition task is composed solely of authentic oracle bone rubbings or fragments, it may be affected by the response preferences of LMMs in Yes-or-No questions.
-In this section, we take a deeper analysis of the Yes-or-No judgment ability of LMMs, that whether these models can get similar accuracy on questions that should be answered with Yes, as those should be replied as No.
+In this section, we take a deeper analysis of the Yes-or-No judgment ability of LMMs, and whether these models can get similar accuracy on questions that should be answered with ‘Yes’, as those should be replied to as ‘No’.
 Specifically, we re-curate a content-mixed test set, which consists of 1,000 OBI images and other images containing ancient texts (such as the inscription on the bronzes, characters engraved on bamboo slips, and characters written on silk).